$2000 total deposit with 2 savings accounts. One pays interest rate of 6%the other at rate of 8% if a total earned interest is $144 how much is each deposit
Answer by mananth(16946)

Part I 6.00% per annum
Part II 8.00% per annum

x+y=2000 ------------------------1
6.00%x+ 8.00%y= $144.00
Multiply by 100
6x + 8 y = $14,400.00 --------2

Multiply (1) by -5
we get

-5x-5y= -10000.00

Add this to (2)

3y= $4,400.00

divide by 3

y=$1,466.67 investment at 8.00%
Balance $533.33 investment at 6.00%